If you roll a single die and count the number of dots on top, what is the sample space of all possible outcomes? Are the outcomes equally likely? 0, 1, 2, 3, 4, 5, 6; equally likely 1, 2, 3, 4, 5, 6; equally likely 0, 1, 2, 3, 4, 5, 6; not equally likely 1, 2, 3, 4, 5, 6; not equally likely (b) Assign probabilities to the outcomes of the sample space of part (a). (Enter your answers as fractions.) Outcome Probability 1 2 3 4 5 6 Do the probabilities add up to 1? Should they add up to 1? Explain. Yes, but they should not because these values do not cover the entire sample space. Yes, because these values cover the entire sample space. No, because these values do not cover the entire sample space. No, but they should because these values cover the entire sample space. (c) What is the probability of getting a number less than 4 on a single throw? (Enter your answer as a fraction.) (d) What is the probability of getting 2 or 3 on a single throw? (Enter your answer as a fraction.)
Answer #1

equally likely 1, 2, 3, 4, 5, 6

a)

outcome Probability
1 1/6
2 1/6
3 1/6
4 1/6
5 1/6
6 1/6

  Yes, because these values cover the entire sample space.

c)

robability of getting a number less than 4 on a single throw =3/6 =1/2

d)

probability of getting 2 or 3 on a single throw =2/6 =1/3